Determine the pH of each aqueous solution

. a. 0.475 M oxalic acid (C2O4H2) and 0.500 M HClO.

b. 0.150 M sodium formate (NaHCO2) and 0.138 M pyridine (C5H5N)
Answer #1

o 0.475 Coon oxalic acid ka, = 5.9x102 kq2 = 6.48105 Соо и acid is 0.500M HUO K9 3.0X10-8 oxalic acid Hoo both are weak acid= -bt b²-490 22 = -5.9x10^2 + 1 (5-4 X10-43 + 4x 2.803x102x1 -5.9x10* £ 34.81x10-4+11.212x102 -5.9 X10 1.115601 = 0.1405 X =bi 0.150M Sodium formate (NaHCO3) and 0.138 M Pyridine kg of HCO3 = 2.088104 Kb J G H N = last 1.5X10-9 H605 is strong base dbi 0.150M Sodium formate (NaHCO3) and 0.138 M Pyridine kg of HCO3 = 2.088104 Kb J G H N = last 1.5X10-9 H605 is strong base dconsider for Pyrindine ² C5 H5 Nht - initial change CGNHS + H₂O 0.138 - tot 0.0056 At 0.0056+2 equilibrium 0.138-X ko= [cghsn